# SPHERE: Semantic-PHysical Engaged REpresentation for 3D Semantic Scene Completion

### v1

@@ 4.3. Quantitative Results
SemanticKITTI: Table 1 presents the comparison results on the SemanticKITTI validation set. It can be observed that our SPHERE approach obtains superior performance of 47.91% IoU for class-agnostic scene completion (SC) and 17.01% mIoU for semantic scene completion (SSC), respectively. Compared to the best comparison method SGN (Mei et al., 2024) with ResNet50 backbone, our SPHERE achieves performance improvements of 1.70% IoU and 1.69% mIoU. Furthermore, compared with CGFormer (Yu et al., 2024), which employs EfficientNetB7 as backbone and Swin-T (Liu et al., 2021) as TPV feature backbone, our SPHERE still improves the SSC results by 1.92% IoU and 0.14% mIoU. The comparison results showcase the effectiveness of our SPHERE in generating more realistic SSC predictions through joint exploitation of semantic and physical information.
SSCBench-KITTI-360: Table 2 demonstrates the comparison results on the SSCBench-KITTI-360 test set, where our SPHERE approach outperforms all comparison methods with the performance of 48.59% SC IoU and 20.56% SSC mIoU, respectively. It is worth noting that Gaussian-based methods (Huang et al., 2024b, a) utilize a total of 38400 Gaussians to represent the scene, resulting in unsatisfying performance due to spatial redundancy. On the other hand, our SPHERE selects focal anchors for Gaussian initialization, achieving superior performance with only 1024 Gaussians around focal regions.
Table 5. Efficiency evaluation against voxel-based method SGN (Mei et al., 2024) and Gaussian-based method GaussianFormer-2 (Huang et al., 2024a).
@@ 4.4. Ablation Studies
Ablation on Network Components. As shown in Table 3, we adopt VoxFormer (Li et al., 2023d) as the baseline method, which achieves $44.15\%$ IoU and $13.5\%$ mIoU. By adding SGI to the baseline, the model performances are improved by $2.08\%$ IoU and $2.02\%$ mIoU respectively, which showcases the effectiveness of leveraging Gaussian representations for improved geometric details. Furthermore, by incorporating PHE, SPHERE is able to model physical-aware contextual details with semantic spherical harmonics and promotes semantic-geometry consistency via focal distribution alignment. Therefore, the SSC performance is further enhanced with improvements of $3.90\%$ IoU and $3.50\%$ mIoU over the baseline method.
@@
Effect of Gaussian numbers. Figure 5 illustrates the effect of the number of Gaussians $K$ in SGI. The best performance is obtained with $K=1024$, much fewer than that in Gaussian-based methods (Huang et al., 2024b, a). This validates SPHERE alleviating the spatial redundancy issue with focal anchors for Gaussian initializations.
@@ 4.5. Qualitative Results
Figure 6 illustrates the visualization results on the SemanticKITTI validation set generated by VoxFormer (Li et al., 2023d), SGN (Mei et al., 2024), and our proposed SPHERE. We highlight the ground truth with cyan boxes as a reference for comparison. The red boxes indicate false predictions of SGN and VoxFormer, while the green boxes highlight improved results from our SPHERE approach. It can be observed that SPHERE generates improved SSC predictions with more accurate semantic predictions and more realistic geometric structures, especially for classes like car, other-vehicle, vegetation, and road. This improvement is attributed to the physical-aware modeling ability of Gaussian representations, and the semantic-geometry consistency promoted by focal distribution alignment.

### v2

@@ 4.3. Quantitative Results
SemanticKITTI: Table 1 presents the comparison results on the SemanticKITTI validation set. It can be observed that our SPHERE approach obtains superior performance of 47.91% IoU for class-agnostic scene completion (SC) and 17.01% mIoU for semantic scene completion (SSC), respectively. Compared to the best comparison method SGN (mei2024camera) with ResNet50 backbone, our SPHERE achieves performance improvements of 1.70% IoU and 1.69% mIoU. Furthermore, compared with CGFormer (yu2024context), which employs EfficientNetB7 as backbone and Swin-T (liu2021swin) as TPV feature backbone, our SPHERE still improves the SSC results by 1.92% IoU and 0.14% mIoU. The comparison results showcase the effectiveness of our SPHERE in generating more realistic SSC predictions through joint exploitation of semantic and physical information.
SSCBench-KITTI-360: Table 2 demonstrates the comparison results on the SSCBench-KITTI-360 test set, where our SPHERE approach outperforms all comparison methods with the performance of 48.59% SC IoU and 20.56% SSC mIoU, respectively. It is worth noting that Gaussian-based methods (huang2024gaussianformer; huang2024probabilistic) utilize a total of 38400 Gaussians to represent the scene, resulting in unsatisfying performance due to spatial redundancy. On the other hand, our SPHERE selects focal anchors for Gaussian initialization, achieving superior performance with only 1024 Gaussians around focal regions.
Table 5. Efficiency evaluation against voxel-based method SGN (mei2024camera) and Gaussian-based method GaussianFormer-2 (huang2024probabilistic).
@@ 4.4. Ablation Studies
Ablation on Network Components. As shown in Table 3, we adopt VoxFormer (li2023voxformer) as the baseline method, which achieves $44.15\%$ IoU and $13.5\%$ mIoU. By adding SGI to the baseline, the model performances are improved by $2.08\%$ IoU and $2.02\%$ mIoU respectively, which showcases the effectiveness of leveraging Gaussian representations for improved geometric details. Furthermore, by incorporating PHE, SPHERE is able to model physical-aware contextual details with semantic spherical harmonics and promotes semantic-geometry consistency via focal distribution alignment. Therefore, the SSC performance is further enhanced with improvements of $3.90\%$ IoU and $3.50\%$ mIoU over the baseline method.
@@
Effect of Gaussian numbers. Figure 5 illustrates the effect of the number of Gaussians $K$ in SGI. The best performance is obtained with $K=1024$, much fewer than that in Gaussian-based methods (huang2024gaussianformer; huang2024probabilistic). This validates SPHERE alleviating the spatial redundancy issue with focal anchors for Gaussian initializations.
@@ 4.5. Qualitative Results
Figure 6 illustrates the visualization results on the SemanticKITTI validation set generated by VoxFormer (li2023voxformer), SGN (mei2024camera), and our proposed SPHERE. We highlight the ground truth with cyan boxes as a reference for comparison. The red boxes indicate false predictions of SGN and VoxFormer, while the green boxes highlight improved results from our SPHERE approach. It can be observed that SPHERE generates improved SSC predictions with more accurate semantic predictions and more realistic geometric structures, especially for classes like car, other-vehicle, vegetation, and road. This improvement is attributed to the physical-aware modeling ability of Gaussian representations, and the semantic-geometry consistency promoted by focal distribution alignment.